1. Find the magnitude of an electron's electric field if the field itself is 50cm away from it.




2. If another electron is placed at this distance, what would be the magnitude of the electrostatic force between the charged particles?




3. Determine the nature of the force.

1
Expert's answer
2022-03-07T10:18:40-0500

1. The electric field due the point charge

"E=k\frac{q}{r^2}"

"E=9*10^9*\frac{1.6*10^{-19}}{0.5^2}=5.76*10^{-9}\:\rm V/m"

2. The force on another electron

"F=qE\\
=1.6*10^{-19}*5.76*10^{-9}=9.2*10^{-28}\:\rm N"

3. Repulsion.